OpenStudy (anonymous):

anyone who can answer this? a point moves on the parabola y^2=8 in such a way that the rate of change of the ordinate is always 5 units/second. how fast is the abscissa changing when the ordinate is 4?

ganeshie8 (ganeshie8):

ordinate : y coordinate abscissa : x coordinate

ganeshie8 (ganeshie8):

could you double check the equation of parabola, is it really `y^2=8` ?

OpenStudy (anonymous):

yes it is y^2=8

OpenStudy (anonymous):

i also think that the questions seems lacking, or something wrong?

ganeshie8 (ganeshie8):

im pretty sure the equation is y^2=8\(\color{red}{x}\)

ganeshie8 (ganeshie8):

y^2 = 8 does not produce a parabola when you graph

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

i also think that way..

ganeshie8 (ganeshie8):

good, lets make it y^2=8x

ganeshie8 (ganeshie8):

we're given \(\dfrac{dy}{dt}=5\)

ganeshie8 (ganeshie8):

\[y^2=8x\] differentiate both sides with respect to \(t\)

OpenStudy (anonymous):

\[2y \frac{ dy }{ dt} = 8\frac{ dx }{ dt }\]

OpenStudy (anonymous):

is that right?

ganeshie8 (ganeshie8):

Yes, plugin the given values

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

\[\frac{ dx }{ dt} = 5\]

OpenStudy (anonymous):

thanks sir

ganeshie8 (ganeshie8):

Yw!